The length of a rectangle is 3 yards less than the twice the width, and the area of the rectangle is 27 yards to the power of two. Find the dimensions of the rectangle.
Answer by Theo(13342)

x equals the length
y equals the width
the length is 3 yards less than twice the width.
x = 2y-3
area is 27^2
length * width = area.
x*y = 27^2
since x = 2y-3, the formula becomes (2y-3)*y = 27^2
simplify to get 2y^2 - 3y = 729
subtract 27^2 from both sides of the equation to get 2y^2 - 3y - 729 = 0
use the quadratic formula to get:
y = 19.856608804285 or y = -18.something.
y can't be negative, so it has to be 19.8566.....
x = 2y-3, so x = 36.71321761
the area of the rectangle is x*y = 19.856608804285 * 36.71321761 = 729.
729 is equal to 27^2, so the requirements of the problem are satisfied.